OpenStudy (anonymous):

Joy drove 120 miles the first day of her trip. She drove 60 mph on the following days. What is the beginning amount (b) in the linear model?

OpenStudy (anonymous):

Same as above, 120 is fixed and thereafter she drives 60 miles each hour. y = 60x +120, where x is in hours. So, b = 120.

OpenStudy (anonymous):

thanks, do you know this one? Andrew drove 55 mph on his trip. Which of the following equations best represents the distance he drove if x stands for the number of hours? @britneyspears

OpenStudy (anonymous):

y = 55x, because rate x time = distance.

OpenStudy (anonymous):

A yoga membership costs $16 and an additional $7 per class. Write a linear equation modeling the cost of a yoga membership?

OpenStudy (anonymous):

Think of it like this: you need to pay $16 for membership regardless of what you do, so that's your minimum payment. You pay another $7 per class, so the amount extra you pay depends on the number of classes you do. Say x is the number of classes you do and y is the total cost. The equation would therefore be: y = 16 + 7x

OpenStudy (anonymous):

You drive a car that holds 15 gallons of gasoline. Your car uses 0.03 gallons per mile. Which equation best represents this information?

OpenStudy (anonymous):

these our the chocies.. y = 0.03x + 15 y = 0.03x - 15 y = -0.03x + 15 none of the above

OpenStudy (anonymous):

Annie paid $12 to enter the state fair. She then paid $2 per ride. This equation represents the amount she spent: y = 2x + 12. Use the equation to predict how much she spent if she rode eight rides @britneyspears